# Renormalon-based resummation for spacelike and timelike QCD quantities whose perturbation expansion has general form

### v1

@@ I Introduction
The study of QCD properties at low-momentum transfers ($\lesssim 1$ GeV) is challenging due to the nonperturbative nature of QCD. There are some efforts to try to unify the high- and low-momentum regimes, and most of them lead to noncontinuous and/or nonoholomorphic transtion between these two regimes. Recently, in BrodECH ; ACECH , certain effective charges with holomorphic connection between the two regimes are supposed to contain information of the theory, but they are unfortunately tied to only one specific observable.
In this work, we present a resummation formalism that evaluates the (leading-twist part of the) QCD observables with a single integral, keeping some of the ideas of the effective charge, but with clear differences. An early version of this resummation was constructed by Neubert Neubert some time ago for the large-$\beta_{0}$ approximation (cf. also NC ). Subsequently, the resummation formalism was extended to any loop-level renmod , leading to an integration involving the (exact) running coupling and an observable-dependent characteristic (weight) function, where the latter can be (approximately) determined by the knowledge of the renormalon structure of the considered observable. These resummations were constructed for the case when the perturbation expansions of the considered observables were series of integer powers of the coupling. In the present work, we extend these resummations to the case of observables whose perturbation expansions have, in general, noninteger powers of the couplings. In principle, such resummations are for the spacelike QCD observables, but here we also extend this formalism to the general case of timelike QCD observables, and present a phenomenological application of this formalism.
@@
In this work, we extend the described method to the resummation of the spacelike observables whose perturbation expansion is ${\cal D}=a(Q^{2})^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$ and timelike quantities with perturbation expansion ${\cal F}(\sigma)=a(\sigma)^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$, where $\nu_{0}$ has in general a noninteger value ($0<\nu_{0}\leq 1$). As an illustration, we will apply this approach to a specific timelike observable ${\cal F}(\sigma)$ (with $\nu_{0}=1/3$), and will perform the renormalon-motivated resummation by using both the pQCD coupling $a$ and a holomorphic coupling $\mathcal{A}$}.
@@ II Resummation for the case of the spacelike observable ${\cal D}(Q^{2})$
Here, $\nu_{0}$ is the index of the first expansion term ($0<\nu_{0}\leq 1$), and $\kappa\equiv\mu^{2}/Q^{2}$ denotes the renormalisation scale parameter which is positive and chosen to be $\kappa\sim 1$. The first expansion is in powers of $a(\kappa Q^{2})$, with the expansion coefficients $d_{n}(\nu_{0};\kappa)$ (note: $d_{0}=1$ by convention). The second expansion is in couplings ${\widetilde{a}}_{\nu_{0}+n}$ which are the noninteger generalisation of the logarithmic derivatives of $a$, with the expansion coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ (note: ${\widetilde{d}}_{0}=1$).111 For the integer $\nu_{0}=1$ we have: ${\widetilde{a}}_{n+1}(Q^{2})\equiv(-1)^{n}(n!\beta_{0}^{n})^{-1}(d/d\ln Q^{2})^{n}a(Q^{2})$. The noninteger powers and the noninteger logarithmic derivatives are related GCAK
@@
$\frac{d}{d\ln\kappa}{\cal D}(Q^{2})$ $=$ ${\widetilde{a}}_{\nu_{0}}(\kappa Q^{2})\left[\frac{d}{d\ln\kappa}{\widetilde{d}}_{0}(\nu_{0},\kappa)\right]+\sum_{n=1}^{\infty}{\widetilde{a}}_{\nu_{0}+n}(\kappa Q^{2})\left[(-\beta_{0})(\nu_{0}+n-1){\widetilde{d}}_{n-1}(\nu_{0};\kappa)+\frac{d}{d\ln\kappa}{\widetilde{d}}_{n}(\nu_{0};\kappa)\right],$ (7)
@@
The above problem is addressed by the following Theorems 2 and 3. We first assume that the renormalon structures are dominated by an infrared (IR) renormalon at $u=p$ ($p>0$), and then comment on the modification for the case of an ultraviolet (UV) renormalon at u=pu=-p.
@@
where ${\widetilde{\cal K}}^{(1)}$} is a constant.131313 We took here $\kappa=1$. We note that then: ${{\widetilde{\cal K}}^{(1)}}(\kappa)=\kappa^{p}{{\widetilde{\cal K}}^{(1)}}$, because $\kappa^{u}=\kappa^{p}[1+{\cal O}((p-u))]$. We point out that in our approach we take into account only the leading (i.e., the most singular) renormalon contributions, i.e., we neglect the relative corrections ${\cal O}((p-u)^{1})$ in the (modified) Borel transforms. According to Eq. (36), these relative corrections then contribute to the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ the relative corrections ${\cal O}(1/|\ln t|)$ (we note that $1/|\ln t|<1$ for $0\leq t<1/e$ ($\approx 0.37$).
@@ III Resummation for timelike observable ${\cal F}(\sigma)$
Theorem 6: Let ${\cal D}(Q^{2})$ be the spacelike observable whose expansions are written in Eqs. (1), and ${\cal F}(\sigma)$ be the corresponding timelike quantity defined in Eq. (40a), and let us assume that the modified Borel of ${\cal D}(Q^{2})$ has the form
@@
As shown in GCAK , this timelike coupling obeys the following differential recursion relation [completely analogous to that of ${\widetilde{a}}_{\nu}$ Eq. (5)]:
@@ IV The use of holomorphic QCD (AQCD) couplings in the formalism
This approach suggests that the construction of an AQCD framework reduces basically to the modelling of the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ at low $\sigma$: $0<\sigma<1\ {\rm GeV}^{2}$. [^16] The most prominent early model of holomorphic coupling has the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ equal to the pQCD $\rho_{a}(\sigma)$ for all $\sigma>0$, including low values $\sigma\lesssim 1\ {\rm GeV}^{2}$. This model can be called the Minimal Analytic framework (MA; named also (F)APT) Shirkov ; SMS ; ShirRev ; FAPT ; BakRev , and thus has $\rho_{{\mathcal{A}}}^{\rm(MA)}(\sigma)=\Theta(\sigma)\rho^{\rm(pt)}(\sigma)$. The difference ${\mathcal{A}}(Q^{2})-a(Q^{2})$ in the MA at large $|Q^{2}|$ is $\sim(\Lambda^{2}_{\rm Lan}/Q^{2})^{1}$, i.e., a relatively slow fall-off when $|Q^{2}|$ increases. A generalisation of (F)APT was proposed in AM1 ; AM2 in the context of the pion form factor. In this regime, it is reasonable to represent the parametrisation of $\rho_{{\mathcal{A}}}(\sigma)$ in terms of a sum of a finite number of Dirac deltas, e.g. two deltas 2dQCD1 ; 2dQCD2 or three deltas 3dAQCD .
@@
For specifying numerically the other seven parameters (${\cal F}_{j}$, $M_{j}^{2}$ for $j=1,2,3$; and $M_{0}^{2}$), we use the following inputs: For large $Q^{2}$ ($|Q^{2}|>1\ {\rm GeV}^{2}$) we require ${\mathcal{A}}(Q^{2})$ to approach the underlying perturbative coupling $a(Q^{2})$ quickly, specifically ${\mathcal{A}}(Q^{2})-a(Q^{2})\sim(\Lambda_{\rm Lan}^{2}/Q^{2})^{5}$, and this gives us four conditions; whereas at low $Q^{2}$ the required behaviour is as suggested by large-volume lattice calculations 3dLVLattice : ${\mathcal{A}}(Q^{2})$ at positive $Q^{2}$ has a local maximum at $Q^{2}\approx 0.135\ {\rm GeV}^{2}$ and for $Q^{2}\to 0$ it behaves as ${\mathcal{A}}(Q^{2})\sim Q^{2}$ ($\to 0$). One of the seven parameters is then still free, which we choose to be the threshold scale of the spectral function, namely $\sigma_{\rm thr.}=M_{1}^{2}$, which is expected to be of the order of the square of the lowest hadronic mass: $M_{1}^{2}\approx m_{\pi}^{2}$ ($\sim 0.15^{2}\ {\rm GeV}^{2}$). Note that the renormalisation scheme we use in this construction is MiniMOM MiniMOM because the large-volume lattice calculations (3dLVLattice ) are performed in this scheme. In addition, we rescale the momenta to the usual $\overline{\rm MS}$-type scheme scales (scaled with $\Lambda^{2}_{\overline{\rm MS}}$), and this is called Lambert MiniMOM (LMM) scheme. We refer for details of construction of ${\mathcal{A}}(Q^{2})$ in such framework to 3dAQCD ; amu3dAQCD .
@@
Furthermore, the timelike analog of this coupling, namely ${\widetilde{\mathfrak{H}}}(\sigma)$ which is the AQCD analog of ${\widetilde{\mathfrak{h}}}_{\nu}$ of Eq. (55), can also be obtained in terms of an integral involving the spectral function $\rho_{{\mathcal{A}}}$ GCAK
@@
${\widetilde{\mathfrak{H}}}_{\nu}(\kappa\sigma)=\frac{\sin(\pi\nu)\beta_{0}^{1-\nu}}{\pi^{2}(1-\nu)}\left\{\pi\sum_{j=1}^{3}\frac{{\cal F}_{j}}{M_{j}^{2}}\Theta(M_{j}^{2}-\sigma)\ln^{1-\nu_{0}}\left(\frac{M_{j}^{2}}{\sigma}\right)+\int_{\Theta(M_{0}^{2}-\sigma)\ln(M_{0}^{2}/\sigma)}^{\infty}dw\;w^{1-\nu_{0}}\rho_{a}(\sigma e^{w})\right\}.$ (77)
@@
Furthermore, due to the Landau cuts of the pQCD coupling $a(Q^{2})$ at low positive $Q^{2}$ ($0\leq Q^{2}\leq\Lambda^{2}_{\rm Lan}$) we have to regularise the pQCD resummations Eqs. (39) is some way to avoid those cuts in the integration of $t$ at low $t$ values. One possibility is a PV-type regularisation

### v2

@@ I Introduction
The study of QCD properties at low-momentum transfers ($\lesssim 1$ GeV) is challenging due to the nonperturbative nature of QCD. There are some efforts to try to unify the high- and low-momentum regimes, and most of them lead to noncontinuous and/or nonholomorphic transition between these two regimes. Recently, in BrodECH ; ACECH , certain effective charges with holomorphic connection between the two regimes are supposed to contain information of the theory, but they are unfortunately tied to only one specific observable.
In this work, we present a resummation formalism that evaluates the (leading-twist part of the) QCD observables with a single integral, keeping some of the ideas of the effective charge, but with clear differences. An early version of this resummation was constructed by Neubert Neubert some time ago for the large-$\beta_{0}$ approximation (cf. also NC ). Subsequently, the resummation formalism was extended to any loop-level renmod , leading to an integration involving the (exact) running coupling and an observable-dependent characteristic (weight) function, where the latter can be (approximately) determined by the knowledge of the renormalon structure of the considered observable. These resummations were constructed for the case when the perturbation expansions of the considered observables were series of integer powers of the coupling. In the present work, we extend these resummations to the case of observables whose perturbation expansions have, in general, noninteger powers of the couplings. In principle, such resummations are for the spacelike QCD observables, but here we also extend this formalism to the general case of timelike QCD observables, and present a phenomenological application of this formalism. Expansions in noninteger powers appear in general for observables whose one-loop associated anomalous dimensions are nonzero.
@@
In this work, we extend the described method to the resummation of the spacelike observables whose perturbation expansion is ${\cal D}=a(Q^{2})^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$ and timelike quantities with perturbation expansion ${\cal F}(\sigma)=a(\sigma)^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$, where $\nu_{0}$ has in general a noninteger value ($0<\nu_{0}\leq 1$). Thus, this formalism allows us to evaluate in a consistent way timelike quantities such as decay rates, heavy quark masses and other quantities related with the heavy meson masses, etc., i.e., quantities especially interesting due to their active experimental development. As an illustration, we will apply this approach to a specific timelike observable ${\cal F}(\sigma)$ (with $\nu_{0}=1/3$), and will perform the renormalon-motivated resummation by using both the pQCD coupling $a$ and a holomorphic coupling $\mathcal{A}$}.
[…]
@@ II Resummation for the case of the spacelike observable ${\cal D}(Q^{2})$
Here, $\nu_{0}$ is the index of the first expansion term ($0<\nu_{0}\leq 1$), and $\kappa\equiv\mu^{2}/Q^{2}$ denotes the renormalisation scale parameter which is positive and chosen to be $\kappa\sim 1$. The first expansion is in powers of $a(\kappa Q^{2})$, with the expansion coefficients $d_{n}(\nu_{0};\kappa)$ (note: $d_{0}=1$ by convention). The second expansion is in couplings ${\widetilde{a}}_{\nu_{0}+n}$ which are the noninteger generalisations of the logarithmic derivatives of $a$, with the expansion coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ (note: ${\widetilde{d}}_{0}=1$). [^1] For the integer $\nu_{0}=1$ we have: ${\widetilde{a}}_{n+1}(Q^{2})\equiv(-1)^{n}(n!\beta_{0}^{n})^{-1}(d/d\ln Q^{2})^{n}a(Q^{2})$. The noninteger powers and the noninteger logarithmic derivatives are related GCAK
@@
$\frac{d}{d\ln\kappa}{\cal D}(Q^{2})=0$ (7)
$=$ ${\widetilde{a}}_{\nu_{0}}(\kappa Q^{2})\left[\frac{d}{d\ln\kappa}{\widetilde{d}}_{0}(\nu_{0},\kappa)\right]+\sum_{n=1}^{\infty}{\widetilde{a}}_{\nu_{0}+n}(\kappa Q^{2})\left[(-\beta_{0})(\nu_{0}+n-1){\widetilde{d}}_{n-1}(\nu_{0};\kappa)+\frac{d}{d\ln\kappa}{\widetilde{d}}_{n}(\nu_{0};\kappa)\right],$
@@
In Appendix D we present an extended discussion on the consistency aspects of this formalism, concentrating on several mathematical and numerical considerations. The Appendix takes into account that the couplings used in practice in this formalism are not pQCD couplings (which have Landau singularities), but are holomorphic couplings (they are free from Landau singularities) discussed in Secs. IV-V, as such couplings allow us to perform the resummations (formally) without ambiguity.
@@
The above problem is addressed by the following Theorems 2 and 3. We first assume that the renormalon structures are dominated by an infrared (IR) renormalon at $u=p$ ($p>0$), and then comment on the modification for the case of an ultraviolet (UV) renormalon at u=pu=-p. The location of the (renormalon) singularities of the (usual) Borel transform of a QCD observable can be obtained by the large-$\beta_{0}$ approach. The strength of these singularities (the power of the poles) is modified, at the level beyond the large-$\beta_{0}$, by the RGE for the strong coupling $\beta$-function, cf. Eq. (6a)] and the RGE for the local operator matrix element (anomalous dimension $\gamma$). The RGEs add to the large-$\beta_{0}$ part of the power a term proportional to the two-loop $\beta$-coefficient $\beta_{1}/\beta_{0}^{2}$, and the term $\gamma_{0}^{(1)}/\beta_{0}$ containing the leading-order anomalous dimension coefficient $\gamma_{0}^{(1)}$, respectively (cf., for example, JPGour ). The general term in a (usual) Borel transform $B(u)$ that represents an IR renormalon at $u=p$ ($p>0$) is schematically $\sim 1/(p-u)^{s_{0}}$ where $s_{0}>0$ is the (noninteger) power of the IR renormalon. A general form of an UV renormalon in $B(u)$ at u=pu=-p (p<0-p<0) is $\sim 1/(p+u)^{s_{0}}$.
@@
where ${\widetilde{\cal K}}^{(1)}$} is a constant. [^13] We took here $\kappa=1$. We note that then: ${{\widetilde{\cal K}}^{(1)}}(\kappa)=\kappa^{p}{{\widetilde{\cal K}}^{(1)}}$, because $\kappa^{u}=\kappa^{p}[1+{\cal O}((p-u))]$. We point out that in our approach we take into account only the leading (i.e., the most singular) renormalon contributions, i.e., we neglect the relative corrections ${\cal O}((p-u)^{1})$ in the (modified) Borel transforms. According to Eq. (36), these relative corrections then contribute to the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ the relative corrections ${\cal O}(1/|\ln t|)$ (we note that $1/|\ln t|<1$ for $0\leq t<1/e$ ($\approx 0.37$).
@@ III Resummation for timelike observable ${\cal F}(\sigma)$
Relations (41) are similar to once-subtracted dispersion relations. We wrote them here for completeness because they are known to be the exact inverse relations of the relations (40).
@@
We will comment on the justification for such an exchange of the order of integration within the holomorphic QCD (AQCD) later on in Sec. V.
@@
Theorem 6: Let ${\cal D}(Q^{2})$ be the spacelike observable whose expansions are written in Eqs. (1), and ${\cal F}(\sigma)$ be the corresponding timelike quantity defined in Eq. (40a), and let us assume that the modified Borel of ${\cal D}(Q^{2})$ has the form [^16] The form (54) for the modified Borel transform ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u)$ was chosen because it corresponds, according to Theorem 2, to the associated general form of the IR renormalon term $\sim 1/(p-u)^{s_{0}}$ (at $u=p>0$ and with power $s_{0}$) in the usual Borel transform $B^{(\nu_{0})}[{\cal D}](u)$, cf. Eqs. (25)-(27).
@@
As shown in GCAK , the timelike coupling Eq. (44) obeys the following differential recursion relation [completely analogous to that of ${\widetilde{a}}_{\nu}$ Eq. (5)]:
@@
Lemma Eq. (68) has the following consequence: Theorem 3 [cf. Eqs. (26) and (28)], that was applied to ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u)$ and ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$, can be applied also to ${\widetilde{B}}^{(\nu_{0})}[{\cal F}](u)$ and ${\widetilde{B}}[{{\cal F}^{(1)}}](u)$:
${\widetilde{B}}^{(\nu_{0})}[{\cal F}](u;\kappa)$ $=$ $\frac{{\widetilde{\cal K}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}}}\left[1+{\cal O}((p-u))\right]\quad\Rightarrow$ (73a)
${\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa)$ $=$ $\frac{{{\widetilde{\cal K}}^{(1)}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}-\nu_{0}+1}}\left[1+{\cal O}((p-u))\right].$ (73b)
This means that, if we know ${\widetilde{B}}^{(\nu_{0})}[{\cal F}](u)$, then ${\widetilde{B}}[{{\cal F}^{(1)}}](u)$ can be obtained. When using subsequently Theorem 5, Eq. (45), the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ is obtained.
Further, we note that Theorem 2 [Eqs. (25)-(27)] is valid also for timelike observables, because in the proof of that Theorem we did not use the spacelike (or timelike) character of the associated observable. Therefore
$B^{(\nu_{0})}[{\cal F}](u;\kappa)$ $=$ $\frac{{\cal K}(\kappa)}{(p-u)^{s_{0}}}\left[1+{\cal O}((p-u))\right]\quad\Rightarrow$ (74a)
${\widetilde{B}}^{(\nu_{0})}[{\cal F}](u;\kappa)$ $=$ $\frac{{\widetilde{\cal K}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}}}\left[1+{\cal O}((p-u))\right],$ (74b)
where the two indices are related
$s_{0}={\widetilde{s}}_{0}+p\frac{\beta_{1}}{\beta_{0}^{2}}.$ (75)
Therefore, we conclude that from the knowledge of the structure of the (usual) Borel transform $B^{(\nu_{0})}[{\cal F}](u)$ of a timelike observable ${\cal F}(\sigma)$, we obtain the structure of the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ of the auxiliary spacelike quantity ${{\cal D}^{(1)}}(Q^{2})$, via the following steps:
$B^{(\nu_{0})}[{\cal F}](u)\mapsto{\widetilde{B}}^{(\nu_{0})}[{\cal F}](u)\mapsto{\widetilde{B}}[{{\cal F}^{(1)}}](u)\mapsto{\widetilde{B}}[{{\cal D}^{(1)}}](u),$ (76)
where the first step is via Eqs. (74), the second step is via Eqs. (73), and the third step is via Theorem 5 [Eq. (45)].
We recall that we usually (or sometimes) know the renormalon structure of the (usual) Borel transform of the considered timelike quantity ${\cal F}(\sigma)$, but in order to resum it we need the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ [cf. Eq. (43)], and to obtain the latter we need to know ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ because $F_{{{\cal D}^{(1)}}}(t)$ is the inverse Mellin transform of ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ [cf. Eq. (16)]. And the structure of ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ is obtained from the structure of $B^{(\nu_{0})}[{\cal F}](u)$ via the steps (76).
@@ IV The use of holomorphic QCD (AQCD) couplings in the formalism
This approach suggests that the construction of an AQCD framework reduces basically to the modelling of the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ at low $\sigma$: $0<\sigma<1\ {\rm GeV}^{2}$. [^17] The most prominent early model of holomorphic coupling has the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ equal to the pQCD $\rho_{a}(\sigma)$ for all $\sigma>0$, including low values $\sigma\lesssim 1\ {\rm GeV}^{2}$. This model can be called the Minimal Analytic framework (MA; named also (F)APT) Shirkov ; SMS ; ShirRev ; FAPT ; SMnu ; BakRev , and thus has $\rho_{{\mathcal{A}}}^{\rm(MA)}(\sigma)=\Theta(\sigma)\rho^{\rm(pt)}(\sigma)$. The difference ${\mathcal{A}}(Q^{2})-a(Q^{2})$ in the MA at large $|Q^{2}|$ is $\sim(\Lambda^{2}_{\rm Lan}/Q^{2})^{1}$, i.e., a relatively slow fall-off when $|Q^{2}|$ increases. A generalisation of (F)APT was proposed in AM1 ; AM2 in the context of the pion form factor. In this regime, it is reasonable to represent the parametrisation of $\rho_{{\mathcal{A}}}(\sigma)$ in terms of a sum of a finite number of Dirac deltas, e.g. two deltas 2dQCD1 ; 2dQCD2 or three deltas 3dAQCD .
@@
For specifying numerically the other seven parameters (${\cal F}_{j}$, $M_{j}^{2}$ for $j=1,2,3$; and $M_{0}^{2}$), we use the following inputs: For large $Q^{2}$ ($|Q^{2}|>1\ {\rm GeV}^{2}$) we require ${\mathcal{A}}(Q^{2})$ to approach the underlying perturbative coupling $a(Q^{2})$ quickly, specifically ${\mathcal{A}}(Q^{2})-a(Q^{2})\sim(\Lambda_{\rm Lan}^{2}/Q^{2})^{5}$, and this gives us four conditions; whereas at low $Q^{2}$ the required behaviour is as suggested by large-volume lattice calculations 3dLVLattice ; Lattb ; Latt2b ; Latt2c : ${\mathcal{A}}(Q^{2})$ at positive $Q^{2}$ has a local maximum at $Q^{2}\approx 0.135\ {\rm GeV}^{2}$ and for $Q^{2}\to 0$ it behaves as ${\mathcal{A}}(Q^{2})\sim Q^{2}$ ($\to 0$). One of the seven parameters is then still free, which we choose to be the threshold scale of the spectral function, namely $\sigma_{\rm thr.}=M_{1}^{2}$, which is expected to be of the order of the square of the lowest hadronic mass: $M_{1}^{2}\approx m_{\pi}^{2}$ ($\sim 0.15^{2}\ {\rm GeV}^{2}$). Note that the renormalisation scheme we use in this construction is MiniMOM MiniMOM because the large-volume lattice calculations (3dLVLattice ; Lattb ) are performed in this scheme. In addition, we rescale the momenta to the usual $\overline{\rm MS}$-type scheme scales (scaled with $\Lambda^{2}_{\overline{\rm MS}}$), and this is called Lambert MiniMOM (LMM) scheme. We refer for details of construction of ${\mathcal{A}}(Q^{2})$ in such framework to 3dAQCD ; amu3dAQCD . [^18] In principle, it is possible to include even more than three Dirac $\delta$-functions in the ansatz for $\rho_{{\mathcal{A}}}(\sigma)$. The number of equations fixing all the parameters then increases, e.g., by requiring very high level of coincidence of ${\mathcal{A}}(Q^{2})$ with the underlying pQCD coupling $a(Q^{2})$ in the high-momentum perturbative regime $|Q^{2}|>1\ {\rm GeV}^{2}$, such as ${\mathcal{A}}(Q^{2})-a(Q^{2})\sim(\Lambda_{\rm Lan}^{2}/Q^{2})^{N_{\rm max}}$ with $N_{\rm max}>5$. However, all these equations are nonlinear, and it becomes increasing hard to solve them numerically as their number increases. Further, in our case of (only) three Dirac delta functions, we can already fulfill the (two) mentioned conditions in the IR regime, and in the UV regime we have $N_{\rm max}=5$ which is sufficient for the coupling ${\mathcal{A}}(Q^{2})$ to practically coincide with the underlying pQCD coupling $a(Q^{2})$ for $Q^{2}>3\ {\rm GeV}^{2}$, cf. Fig. 1.
@@
Furthermore, the timelike analog of this coupling, namely ${\widetilde{\mathfrak{H}}}_{\nu}(\sigma)$ which is the AQCD analog of ${\widetilde{\mathfrak{h}}}_{\nu}$ of Eq. (44), can also be obtained in terms of an integral involving the spectral function $\rho_{{\mathcal{A}}}$ GCAK
@@
${\widetilde{\mathfrak{H}}}_{\nu}(\kappa\sigma)=\frac{\sin(\pi\nu)\beta_{0}^{1-\nu}}{\pi^{2}(1-\nu)}\left\{\pi\sum_{j=1}^{3}\frac{{\cal F}_{j}}{M_{j}^{2}}\Theta(M_{j}^{2}-\sigma)\ln^{1-\nu}\left(\frac{M_{j}^{2}}{\sigma}\right)+\int_{\Theta(M_{0}^{2}-\sigma)\ln(M_{0}^{2}/\sigma)}^{\infty}dw\;w^{1-\nu}\rho_{a}(\sigma e^{w})\right\}.$ (81)
The differential recursion relation (56) remains valid in AQCD GCAK , i.e., with the substitutions $\widetilde{\mathfrak{h}}$$\mapsto{\widetilde{\mathfrak{H}}}$.
@@
Furthermore, due to the Landau cuts of the pQCD coupling $a(Q^{2})$ at low positive $Q^{2}$ ($0\leq Q^{2}\leq\Lambda^{2}_{\rm Lan}$) we have to regularise the pQCD resummations Eqs. (39) in some way to avoid those cuts in the integration of $t$ at low $t$ values. One possibility is a PV-type regularisation